how long would it take a money to double if it is invested at 7 % compounded daily? 8.3 % compounded continuously ??
Answer by nerdybill(7384)

how long would it take a money to double if it is invested at 7 % compounded daily? 8.3 % compounded continuously ??
.
Continuous compound equation:
A = Poe^(rt)
where
A is amount at time t
Po is initial amount
r is rate or interest
t is time
.
Let x = initial investment
2x = xe^(.083t)
2 = e^(.083t)
ln(2) = .083t
ln(2)/.083 = t
8.35 years = t